OpenStudy (anonymous):

Solve 3x – 5y = 15 for y.

OpenStudy (anonymous):

3x - 5y = 15 +5y +5y --------------- 3x = 15 + 5y 3x = 15 + 5y -15 -15 -------------- 3x - 15 = 5y 5y = (3x - 15) /5 /5 ------------- y = 3/5x - 3